OpenStudy (anonymous):

Bank of America customers select their own three- digit personal identification number (PIN) for use at ATMs. a. Think of this as an experiment and list four possible outcomes. b. What is the probability Mr. Jones and Mrs. Smith select the same PIN? c. Which concept of probability did you use to answer (b)?

OpenStudy (amistre64):

a, your smart enough to do for yourself. b, hmm there are at most 999 + 000 combos; soo 1000 altogether 1/1000 + 1/1000 = 2/1000 = 1/500

OpenStudy (amistre64):

c, you used the concept of openstudy lol

OpenStudy (amistre64):

or is that multiplication?

OpenStudy (unklerhaukus):

concept of randomness

OpenStudy (amistre64):

i think b might be 1/1000000 youd have to dbl chk my idiocy on that one

OpenStudy (anonymous):

Okay thank you. :)

OpenStudy (unklerhaukus):

probability is 1 in 1000, , there is only 1 degree of freedom

OpenStudy (anonymous):

Oh I see.